OpenStudy (loser66):

Let K be one of Q, R, C (or any field). Prove that a polynomial f(x) in K[x] of degree 3 is reducible if and only if f(x) has a root in K. I need help on the backward proof. I mean "If f has a root, then f is reducible". Please

OpenStudy (loser66):

I got the forward one. "if f is reducible, f has a root."

ganeshie8 (ganeshie8):

Is it like, if f has a root, then it can be expressed as a product of two factors ?

OpenStudy (loser66):

Yes, I think so

OpenStudy (loser66):

not just 2, can be 3

ganeshie8 (ganeshie8):

i wont be useful here

OpenStudy (loser66):

Thanks anyway. It's better than no one come. :)

ganeshie8 (ganeshie8):

hope somebody assists sooner, good luck! @SithsAndGiggles

OpenStudy (loser66):

@kirbykirby

OpenStudy (loser66):

oh, nvm, I got it. hihihi If \(\alpha \) is a root of f then \((x-\alpha) |f(x)\) ) or f(x) = \((x-\alpha) *q(x) That shows f(x) is reducible